# SeaS: Few-shot Industrial Anomaly Image Generation with Separation and Sharing Fine-tuning

### v1

@@ A.6 More qualitative and quantitative anomaly image generation results
Category DRAEM DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
bottle 96.40 98.56 94.12 96.98 98.76 95.35 100.00 100.00 100.00 100.00 100.00 100.00
cable 70.91 78.06 71.52 90.98 94.21 88.14 99.52 99.55 97.71 92.05 94.95 88.70
capsule 85.80 95.77 87.80 86.32 95.99 88.46 83.25 94.62 89.44 93.80 98.19 93.42
carpet 81.28 92.34 84.67 88.02 95.33 87.60 86.00 93.42 87.22 97.98 99.22 96.67
grid 97.14 98.63 92.86 85.48 92.61 85.71 93.69 97.08 91.14 96.79 98.76 96.10
hazelnut 85.73 89.00 81.72 99.90 99.91 98.97 98.28 98.60 95.83 100.00 100.00 100.00
leather 98.81 99.28 98.44 95.93 98.15 93.65 99.90 99.95 99.20 100.00 100.00 100.00
metal_nut 94.67 98.39 93.44 96.16 98.57 96.18 99.01 99.65 98.46 98.58 99.54 97.64
pill 66.39 89.29 88.07 82.85 94.40 92.00 94.15 98.42 94.47 98.16 99.50 96.84
screw 73.92 85.38 82.72 82.60 92.15 82.22 81.54 91.32 82.05 87.83 94.39 85.54
tile 95.80 97.49 93.33 98.94 99.43 96.55 98.25 99.13 95.65 99.36 99.69 99.12
toothbrush 83.33 90.29 84.44 77.08 87.68 80.95 100.00 100.00 100.00 87.92 94.08 87.80
transistor 90.06 89.06 81.48 88.04 85.06 77.78 97.38 96.57 92.86 98.10 96.90 94.55
wood 99.50 99.78 97.62 99.87 99.94 98.82 97.24 98.70 96.47 100.00 100.00 100.00
zipper 99.39 99.77 97.56 97.07 98.78 96.25 99.01 99.71 99.39 100.00 100.00 100.00
Average 87.94 93.41 88.65 91.08 95.40 90.58 95.15 97.78 94.66 96.70 98.35 95.76
A.7 More details of the segmentation models
As mentioned in the experiment part, we choose three segmentation models (BiSeNet V2 (Yu et al., 2021), UPerNet (Xiao et al., 2018), LFD (Zhou et al., 2024a)) to verify the validity of the generated image-mask pairs on the downstream anomaly segmentation as well as detection tasks. For BiSeNet V2 and UPerNet, we generally follow the implementation provided by MMsegmentation. For LFD, we also use the official implementation.
In training segmentation models for downstream tasks, we adopt a training strategy of training a unified segmentation model for all classes of products, rather than training separate segmentation models for each class. Experimental results are shown in Tab. 19, which indicate that the performance of the unified segmentation model surpasses that of multiple individual segmentation models.
Table 19: Ablation on the training strategy of segmentation models.
Models Multiple Models Unified Model
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
BiSeNet V2 96.00 67.68 65.87 54.11 97.21 69.21 66.37 55.28
UPerNet 96.77 73.88 70.49 60.37 97.87 74.42 70.70 61.24
LFD 93.02 72.97 71.56 55.88 98.09 77.15 72.52 56.47
Average 95.26 71.51 69.31 56.79 97.72 73.59 69.86 57.66
A.8 More ablation studies
Ablation on the Unbalanced Abnormal Text Prompt design
we are able to provide high-fidelity and diverse images for downstream anomaly segmentation tasks, resulting in the best performance in segmentation metrics.
Ablation on the Separation and Sharing Fine-tuning loss
where $A^{l}_{\text{ob}}\in\mathbb{R}^{r\times r\times 1}$ is the cross-attention map corresponding to the normal token <ob>. As shown in Tab. 21, the experimental results demonstrate that, our adopted loss design achieves the best performance in downstream segmentation tasks.
Table 20: Ablation on the Unbalanced Abnormal Text Prompt design.
Prompt AUROC AP $F_{1}$-max IoU
$N^{\prime}$ = 1, $N$ = 1 96.48 63.69 62.50 52.02
$N^{\prime}$ = 1, $N$ = 4 (Ours) 97.21 69.21 66.37 55.28
$N^{\prime}$ = 4, $N$ = 4 96.55 66.28 63.95 54.07
Table 21: Ablation on the Separation and Sharing Fine-tuning loss.
Loss AUROC AP $F_{1}$-max IoU
w/o ST 96.44 67.73 65.23 54.99
with AT 96.42 63.99 62.43 53.36
Ours 97.21 69.21 66.37 55.28
Ablation on the minimum size requirement for training images
In the few-shot setting, for a fair comparison, we follow the common setting in DFMGAN (Duan et al., 2023) and AnomalyDiffusion (Hu et al., 2024), i.e., using one-third abnormal image-mask pairs for each anomaly type in training. In this setting, the minimum number of abnormal training images is 2. Once we adopt a 3-shot setting, we need to reorganize the test set. To ensure that the test set is not reorganized for fair comparison, we take 1-shot and 2-shot settings for all anomaly types during training, i.e., $H=1$ and $H=2$, where $H$ is the image number. The results are shown in Tab. 22 and Fig. 15. Observably, the models trained by 1-shot and 2-shot settings still generate anomaly images with decent diversity and authenticity.
Size IS IC-L
$H=1$ 1.790 0.311
$H=2$ 1.794 0.314
$H=\frac{1}{3}\times H_{0}$ 1.876 0.339
Figure 15: Visualization of the ablation study on the minimum size requirement for training images. In the figure, the first row is for generated images, the second row is for generated masks.
Ablation on the training strategy of SeaS
Table 23: Ablation on training strategy of SeaS.
Strategy IS IC-L
Abnormal-Normal 1.53 0.28
Normal-Abnormal 1.70 0.32
Abnormal&Normal (Ours) 1.88 0.34
In Decoupled Anomaly Alignment (DA) loss, we leverage cross-attention maps from various layers of the U-Net encoder. Specifically, we investigate the impact of integrating different cross-attention maps, denoted as $A^{1}\in\mathbb{R}^{64\times 64}$, $A^{2}\in\mathbb{R}^{32\times 32}$, $A^{3}\in\mathbb{R}^{16\times 16}$ and $A^{4}\in\mathbb{R}^{8\times 8}$. These correspond to the cross-attention maps of the “down-1”, “down-2”, “down-3”, and “down-4” layers of the encoder in U-Net respectively. As shown in Tab. 25, the experimental results demonstrate that, employing a combination of $\{A^{2},A^{3}\}$ for DA loss, achieves the best performance in downstream segmentation tasks.
Ablation on the features for Coarse Feature Extraction
In the coarse feature extraction process, we extract coarse but highly-discriminative features for anomalies from U-Net decoder. Specifically, we investigate the impact of integrating different features, denoted as $F_{1}\in\mathbb{R}^{16\times 16\times 1280}$, $F_{2}\in\mathbb{R}^{32\times 32\times 1280}$, $F_{3}\in\mathbb{R}^{64\times 64\times 640}$ and $F_{4}\in\mathbb{R}^{64\times 64\times 320}$. These correspond to the output feature “up-1”, “up-2”, “up-3”, and “up-4” layers of the encoder in U-Net respectively. As shown in Tab. 25, the experimental results demonstrate that, employing a combination of $\{F_{2},F_{3}\}$ for coarse feature extraction, achieves the best performance in downstream segmentation task.
$A^{l}$ AUROC AP $F_{1}$-max IoU
$l$ = 1, 2, 3 96.42 68.92 66.24 54.52
$l$ = 2, 3, 4 95.71 64.51 62.33 52.46
$l$ = 2, 3 (Ours) 97.21 69.21 66.37 55.28
Table 25: Ablation on the features for Coarse Feature Extraction.
$F_{y}$ AUROC AP $F_{1}$-max IoU
$y$ = 1, 2, 3 94.35 63.58 60.54 52.36
$y$ = 2, 3, 4 96.93 67.42 64.26 55.31
$y$ = 2, 3 (Ours) 97.21 69.21 66.37 55.28
Ablation on the features of VAE for Refined Mask Prediction
In the Refined Mask Prediction, we combine the high-resolution features of VAE decoder features with discriminative features from U-Net, to generate accurately aligned anomaly image-mask pairs. In addition, we can also use the VAE encoder features as high-resolution features. As shown in Tab. 26, the experimental results show that, using VAE decoder features achieves better performance in downstream segmentation tasks.
Table 26: Ablation on the features of VAE for Refined Mask Prediction.
$F^{res}$ AUROC AP $F_{1}$-max IoU
VAE encoder 96.14 66.26 63.48 54.99
VAE decoder 97.21 69.21 66.37 55.28
Ablation on the normal image supervision for Refined Mask Prediction
In the Refined Mask Prediction branch, we predict masks for normal images as the supervision for the mask prediction. We conduct two sets of experiments: (a) We remove the second and the fourth term in the loss for RMP, i.e., the normal image supervision (short for NIA in Tab. 27); (b) We use the complete form in RMP branch loss, i.e., we use the normal image for supervision, as in Eq. equation 10:
$\mathcal{L}_{\mathcal{M}}=\mathcal{F}(\hat{M}_{\text{df}},\mathbf{M}_{\text{df}})+\mathcal{F}(\hat{M}_{\text{ob}},\mathbf{M}_{\text{ob}})+\mathcal{F}(\hat{M}_{\text{df}}^{\prime},\mathbf{M}_{\text{df}}^{\prime})+\mathcal{F}(\hat{M}_{\text{ob}}^{\prime},\mathbf{M}_{\text{ob}}^{\prime})$ (10)
As shown in Tab. 27, the experimental results show that, using normal images for supervision achieves better performance in downstream segmentation tasks. We also provide further qualitative results of the effect of normal image supervision (short for NIA in Fig. 16) on MVTec AD.
$F^{res}$ AUROC AP $F_{1}$-max IoU
w/o NIA 96.20 66.03 64.09 53.97
with NIA (Ours) 97.21 69.21 66.37 55.28
Ablation on the Mask Refinement Module
In the Refined Mask Prediction branch, the Mask Refinement Module (MRM) is utilized to generate refined masks. We devise different structures for MRM, as shown in Fig. 17, including Case a): those without conv blocks, Case b): with one conv blocks, and Case c): with chained conv blocks. As shown in Fig. 18, we find that using the conv blocks in Case b), which consists of two 1 $\times$ 1 convolutions and one 3 $\times$ 3 convolution, helps the model learn the features of the defect area more accurately, rather than focusing on the background area for using one convolution alone in Case a). Based on this observation, we further designed a chained conv blocks structure in Case c), and the acquired features better reflect the defect area. This one-level-by-one level of residual learning helps the model achieve better residual correction results for the defect area features. As shown in Tab. 28 in the Appendix, Case c) improves the performance by + 0.28% on AUROC, + 2.29% on AP and + 2.29% on $F_{1}$-max, + 0.32% on IoU compared with Case b). We substantiate the superiority of the MRM structures that we design, through the results of downstream segmentation experiments.
Figure 18: Visualization of the MRM module intermediate results. The top is for MRM structure diagram, and the bottom is sequentially for input image, feature maps of MRM intermediate process and the predicted mask.
Table 28: Ablation on the Mask Refinement Module.
Model AUROC AP $F_{1}$-max IoU
with MRM (a) 96.75 68.18 64.96 55.51
with MRM (b) 96.93 66.92 64.08 54.96
with MRM (c) 97.21 69.21 66.37 55.28
Ablation on the threshold for mask binarization
In the Refined Mask Prediction branch, we take the threshold $\tau$ for the second channel of refined anomaly masks $\hat{M}_{\text{df}}^{\prime}$ to segment the final anomaly mask. We train segmentation models using anomaly masks with $\tau$ settings ranging from 0.1 to 0.5. As shown in Tab. 29, results indicate that setting $\tau$ = 0.2 yields the best model performance.
Table 29: Ablation on the threshold for mask binarization.
threshold AUROC AP $F_{1}$-max IoU
$\tau$ = 0.1 97.56 65.33 63.38 52.40
$\tau$ = 0.2 (Ours) 97.21 69.21 66.37 55.28
$\tau$ = 0.3 97.20 66.92 64.35 54.68
$\tau$ = 0.4 95.31 63.55 61.97 53.03
$\tau$ = 0.5 94.11 60.85 59.92 50.87
A.9 More qualitative comparison results of segmentation models trained on image-mask pairs generated by different anomaly generation methods
Figure 19: Comparison results with the anomaly segmentation models on MVTec AD. In the figure, from top to bottom are the results for bottle, cable, capsule and carpet categories.
Figure 20: Comparison of the anomaly segmentation results on MVTec AD. In the figure, from top to bottom are the results for grid, hazelnut, leather and metal_nut categories.
Figure 21: Comparison of the anomaly segmentation results on MVTec AD. In the figure, from top to bottom are the results for pill, screw, tile and toothbrush categories.
Figure 22: Comparison of the anomaly segmentation results on MVTec AD. In the figure, from top to bottom are the results for transistor, wood and zipper categories.
A.10 More qualitative comparison results of different segmentation models trained on image-mask pairs generated by SeaS
In this section, we provide further qualitative results with different segmentation models on the MVTec AD dataset. We choose three models with different parameter quantity scopes (BiSeNet V2 (Yu et al., 2021): 3.341M, UPerNet (Xiao et al., 2018): 64.042M, LFD (Zhou et al., 2024a): 0.936M). We report the segmentation results of SeaS for varying types of anomalies in each category. Results are from Fig. 23 to Fig. 26.
Figure 23: Qualitative comparison results with the segmentation models on MVTec AD. In the figure, from top to bottom are the results for transistor, wood, and zipper categories.
Figure 24: Qualitative comparison results with the segmentation models on MVTec AD. In the figure, from top to bottom are the results for bottle, cable, capsule and carpet categories.
Figure 25: Qualitative comparison results with the segmentation models on MVTec AD. In the figure, from top to bottom are the results for grid, hazelnut, leather and metal_nut categories.
Figure 26: Qualitative comparison results with the segmentation models on MVTec AD. In the figure, from top to bottom are the results for pill, screw, tile and toothbrush categories.

### v2

@@ A.6 More qualitative and quantitative anomaly image generation results
Category DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
bottle 96.98 98.76 95.35 100.00 100.00 100.00 100.00 100.00 100.00
cable 90.98 94.21 88.14 99.52 99.55 97.71 92.05 94.95 88.70
capsule 86.32 95.99 88.46 83.25 94.62 89.44 93.80 98.19 93.42
carpet 88.02 95.33 87.60 86.00 93.42 87.22 97.98 99.22 96.67
grid 85.48 92.61 85.71 93.69 97.08 91.14 96.79 98.76 96.10
hazelnut 99.90 99.91 98.97 98.28 98.60 95.83 100.00 100.00 100.00
leather 95.93 98.15 93.65 99.90 99.95 99.20 100.00 100.00 100.00
metal_nut 96.16 98.57 96.18 99.01 99.65 98.46 98.58 99.54 97.64
pill 82.85 94.40 92.00 94.15 98.42 94.47 98.16 99.50 96.84
screw 82.60 92.15 82.22 81.54 91.32 82.05 87.83 94.39 85.54
tile 98.94 99.43 96.55 98.25 99.13 95.65 99.36 99.69 99.12
toothbrush 77.08 87.68 80.95 100.00 100.00 100.00 87.92 94.08 87.80
transistor 88.04 85.06 77.78 97.38 96.57 92.86 98.10 96.90 94.55
wood 99.87 99.94 98.82 97.24 98.70 96.47 100.00 100.00 100.00
zipper 97.07 98.78 96.25 99.01 99.71 99.39 100.00 100.00 100.00
Average 91.08 95.40 90.58 95.15 97.78 94.66 96.70 98.35 95.76
A.8 More qualitative comparison results of different supervised segmentation models trained on image-mask pairs generated by SeaS
In this section, we provide further qualitative results with different supervised segmentation models on the MVTec AD and MVTec 3D AD datasets. We choose three models with different parameter quantity scopes (BiSeNet V2 [40]: 3.341M, UPerNet [38]: 64.042M, LFD [45]: 0.936M). We report the segmentation results of SeaS for varying types of anomalies in each category. Results are from Fig. 23 to Fig. 27.
Figure 23: Qualitative comparison results with the supervised segmentation models on MVTec AD. In the figure, from top to bottom are the results for bottle, cable, capsule and carpet categories.
Figure 24: Qualitative comparison results with the supervised segmentation models on MVTec AD. In the figure, from top to bottom are the results for grid, hazelnut, leather and metal_nut categories.
Figure 25: Qualitative comparison results with the supervised segmentation models on MVTec AD. In the figure, from top to bottom are the results for pill, screw, tile and toothbrush categories.
Figure 26: Qualitative comparison results with the supervised segmentation models on MVTec AD. In the figure, from top to bottom are the results for transistor, wood, and zipper categories.
Figure 27: Qualitative comparison results with the supervised anomaly segmentation models on MVTec 3D AD. In the figure, from top to bottom are the results for cable_gland, carrot, dowel, foam and peach categories.
A.9 Comparison with the Textual Inversion
We conduct the experiment of only using the Textual Inversion (TI) [12] method to learn the product, and the generated images are shown in Fig. 28. The TI method struggles to generate images similar to the real product due to the limited number of learnable parameters. In contrast, for the AIG method, the products satisfy global consistency with minor variations in local details, while the anomalies hold randomness, so the generated products should be globally consistent with the real products. Therefore, unlike the AG method AnomalyDiffusion [17], where the TI method alone is sufficient to meet the anomaly generation needs, we fine-tune the U-Net to ensure the global consistency of the generated products.
Figure 28: Qualitative comparison on the generation results with Textual Inversion.
A.10 More experiments on lighting conditions
We choose one defect class from peach, a product in the MVTec3D dataset, that has significant variations in lighting conditions and backgrounds, to conduct experiments. Images with strong lighting conditions depict the top side of the peach, whereas those with weak lighting conditions show the bottom side. Consequently, the background in the images, whether the top or bottom of the peach, also differs. We selected three training sets with different lighting conditions for experiments: 1) only images from the top side with strong lighting condition, 2) only images from the bottom side with weak lighting condition, 3) half of the images from the top side with strong lighting condition, and a half from the bottom side with weak lighting condition. The generated images of different settings are shown in Fig. 29. It can be seen that SeaS is robust against lighting conditions and background variations.
Figure 29: Visualization of the generation results on MVTec3D AD on different lighting conditions and backgrounds. In the figure, the first row is for the training images and the second row is for the generated images.
A.11 More results on generation of small defects.
SeaS is capable of preserving fine-grained details in small-scale anomalies, as shown in Fig. 30. However, generating extremely subtle anomalies may be challenging due to the limited resolution of the latent space. We will explore this point in our future work.
Figure 30: Generation results of small-scale anomalies.
A.12 More analysis on generation of unseen anomaly types.
SeaS can generate diverse unseen anomalies within known anomaly types as analyzed in Appendix A.2. However, generating truly unseen anomaly types remains challenging. The t-SNE visualizations in Fig. 31 show that different types of anomalies of the same product form compact clusters. Intra-cluster variation is achievable, but cross-cluster generalization is limited by the lack of prior knowledge. We believe that generalizing to unseen anomaly types is important and will explore this in future work.
Figure 31: T-SNE visualization of different anomaly types of the same product in real and generated data.
A.13 More experiments on comparison with DRAEM.
As shown in Tab. 28, training DRAEM[41] with the same anomaly images used in SeaS leads to better results than using only anomaly-free images. However, DRAEM + SeaS achieves further improvements, demonstrating that the gain is not only from real anomalies but also from the diverse and realistic anomalies generated by SeaS.
Table 28: Comparison on combining generated anomalies with synthesis-based anomaly detection method across multiple datasets.
Segmentation Models MVTec AD VisA MVTec 3D AD
Image-level Pixel-level Image-level Pixel-level Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU
DRAEM 98.00 98.45 96.34 97.90 67.89 66.04 60.30 86.28 85.30 81.66 92.92 17.15 22.95 13.57 79.16 90.90 89.78 86.73 14.02 17.00 12.42
DRAEM + training data 97.43 98.84 97.84 96.41 74.42 71.86 59.84 83.74 86.00 82.75 94.63 39.22 43.06 29.02 73.86 88.46 86.69 82.43 19.36 25.05 17.01
DRAEM + SeaS 98.64 99.40 97.89 98.11 76.55 72.70 58.87 88.12 87.04 83.04 98.45 49.05 48.62 35.00 85.45 93.58 90.85 95.43 20.09 26.10 17.07